# RTLCoder: Fully Open-Source and Efficient LLM-Assisted RTL Code Generation Technique

### v4

@@ VI Conclusion
This work proposes a new LLM solution named RTLCoder for RTL code generation, achieving state-of-the-art performance in non-commercial solutions and outperforming GPT-3.5. We contribute a new data generation flow and a complete dataset with over 27 thousand instruction-answer samples, addressing the serious data availability problem in hardware-design-related tasks. Also, we contribute a new training scheme based on design quality scoring. It greatly boosts the model’s performance. Importantly, RTLCoder has been fully open-sourced. RTLCoder’s lightweight property and low hardware barrier allow anyone to easily replicate and further improve based on our existing solution. We expect more brilliant LLM-based solutions in this agile hardware design direction.

### v5

@@ V Conclusion
This work presents a fully open-sourced LLM solution named RTLCoder for RTL code generation, achieving state-of-the-art performance in non-commercial solutions and outperforming GPT-3.5. We contribute a new data generation flow and a complete dataset with over 27 thousand labeled samples, addressing the serious data availability problem in hardware-design-related tasks. Also, we contribute a new training scheme based on design quality scoring. It greatly boosts the model performance. RTLCoder’s lightweight property and low hardware barrier allow anyone to easily replicate and further improve based on our existing solution.
VI Acknowledgement
This work is partially supported by the Hong Kong Research Grants Council (RGC) ECS Grant 26208723.